Problem 10-10

Grand Banks Mining Inc. plans a project to strip-mine a wilderness area. Setting up operations and initial digging will cost $5 million. The first year's operations are expected to be slow and to net a positive cash flow of only $500,000. Then there will be four years of $2 million cash flows after which the ore will run out. Closing the mine and restoring the environment in the sixth year will cost $1 million.

  1. Calculate the project's NPV at a cost of capital of 11%. Enter your answers in dollars and not in millions of dollars. Use a minus sign to indicate a negative NPV. Round PVF values in intermediate calculations to four decimal places. Do not round any other intermediate calculations. Round your answer to the nearest dollar.
    $
  2. Calculate the project's IRR. Do not round intermediate calculations. Round your answer to nearest whole percentage.
    %

10-17

Callaway Associates, Inc. is considering the following mutually exclusive projects. Callaway's Cost of capital is 11%.

Year Project A Project B
0 ($80,000) ($80,000)
1 $44,000 $65,000
2 $34,000 $30,000
3 $14,000 $0
4 $14,000 $5,000
  1. Calculate each project's NPV and IRR. Round the answers to two decimal places.
    Project A Project B
    NPV: $ $
    IRR: % %

  2. Which project should be undertaken? Project A or Project B

10-2

A project has the following cash flows.

C0 C1 C2 C3
($700) $200 $500 $244
  1. What is the project's payback period? Round the answer to two decimal places.
        year(s)
  2. Calculate the project's NPV at 12%. Do not round intermediate calculations. Round PVF and PVFA values in intermediate calculations to four decimal places. Round the answer to two decimal places.
    $  
  3. Calculate the project's PI at 12%. Do not round intermediate calculations. Round PVF and PVFA values in intermediate calculations to four decimal places. Round the answer to two decimal places.
Answer #1

10 -10)

a)

  

NPV = $505,802 (only final answer is rounded to whole dollar amount)

b)

IRR = 15%

10 -17)

a)

Project A:

NPV:

IRR:

Project -B:

NPV:

IRR:

b)

Project A has high NPV where as Project B has high IRR.based on NPV Project A should be selected.

10 -2)

a) pay back period is how fast a project recovers its initial cash investment

here initial cash outflow = 700

cumulative cash flows of year 1 and 2 = 200 + 500 = $700

so pay back period = 2years

b)

NPV = $50.84

c)

profitability index = present value of future cash flows / initial cash outflow

present value of cash flows = (200 / 1.12) + 500 / (1.12^2) + 244 / (1.12^3)

= 750.8427

PI = 750.8427 / 700

= 1.07